### v1

# Algorithmic Collusion And The Minimum Price Markov Game

Abstract
This paper introduces the Minimum Price Markov Game (MPMG), a dynamic variant of the Prisoner’s Dilemma. The MPMG serves as a theoretical model and reasonable approximation of real-world first-price sealed-bid public auctions that follow the minimum price rule. The goal is to provide researchers and practitioners with a framework to study market fairness and regulation in both digitized and non-digitized public procurement processes, amidst growing concerns about algorithmic collusion in online markets. We demonstrate, using multi-agent reinforcement learning-driven artificial agents, that algorithmic tacit coordination is difficult to achieve in the MPMG when cooperation is not explicitly engineered. Paradoxically, our results highlight the robustness of the minimum price rule in an auction environment, but also show that it is not impervious to full-scale algorithmic collusion. These findings contribute to the ongoing debates about algorithmic pricing and its implications.
Igor Sadoune1,2,∗, Marcelin Joanis1,2,‡, Andrea Lodi3,‡
1Department of Mathematics and Industrial Engineering, Polytechnique Montreal, Montreal, QC, Canada
2CIRANO, Montreal, QC, Canada
3Jacobs Technion-Cornell Institute, Cornell Tech and Technion - IIT, New York, NY, USA
Corresponding author: igor.sadoune@polymtl.ca.
These authors contributed equally to this work: marcelin.joanis@polymtl.ca, andrea.lodi@cornell.edu.
1 Introduction
Concerns about algorithmic pricing have grown as technological advancements in machine learning, widespread data availability, and the digitization of the economy have created fertile grounds for such practices [1, 2, 3, 4]. Algorithmic pricing can potentially lead to algorithmic collusion, as observed in digital markets such as online retail [5, 6] and electricity supply markets [7, 8, 9, 10, 11]. These sectors demonstrate the potential for artificial decision-makers to learn collusive behaviors through dynamic pricing algorithms and optimal bidding strategies. This echoes the concerns about algorithmic collusion in public procurement, which are increasingly relevant, especially as digital platforms and automated decision-making tools become more prevalent in these settings. Since public procurement is a crucial mechanism for collective welfare, the vulnerability of this sector to collusive outcomes is particularly alarming. For example, it has been shown that bidding patterns were incompatible with a competitive equilibrium in Ukrainian E-procurement data [12]. Given these developments, questions arise. Is public procurement directly threatened by algorithmic pricing? And if so, under which conditions?
In contrast to digital markets, public procurement in most Western economies typically adheres to first-price sealed-bid auction rules, and more specifically, to the minimum price rule that determines the winner by selecting the lowest bid. This rule is supported by a longstanding body of literature that demonstrates the theoretical efficiency of first-price auctions and, by extension, the efficacy of the minimum price rule in maximizing collective welfare [13, 14, 15]. Public procurement processes, generally not digitized, have made more data available to the public due to recent transparency laws, a response to a long history of traditional cartel collusion [16, 17, 18]. This initiative has been effective as it has helped regulators and researchers detect collusion in public markets [19, 20]. However, the data can also be used to train artificial decision-makers to provide optimal bidding strategies. The potential for these strategies to lead to algorithmic collusion, whether accidental or deliberately learned, raises concerns for legal authorities because such collusion would not trigger antitrust laws [21]. Indeed, accidental algorithmic collusion stems from profit-maximization motives, and proving the deliberate formation of algorithmic coordination remains challenging.
In this paper, we introduce the Minimum Price Markov Game (MPMG), a model designed to provide a framework for studying emergent behavior in the context of minimum price auctions, which are commonly used in real-world public procurement. We draw on Robert Axelrod’s concept of "the emergence of cooperation among egoists" [22], exploring tacit coordination among isolated agents in the MPMG. Departing from the traditional game-theoretic Bayesian framework of first-price auctions, we consider a game of complete information based on the assumption of data availability. Indeed, public auction data typically includes the features of auctions and firms for past and ongoing contracts, as well as the associated bids [23]. This allows us to base the MPMG on the Prisoner’s Dilemma (PD), leveraging a structure that captures the essence of public procurement as a social dilemma—competitive by law but cooperative by nature. Indeed, the PD is well known for encapsulating the paradox of rationality central to our discussion on the study of emergent coordination among isolated agents in public markets. Furthermore, we test the MPMG using various state-of-the-art Multi-Agent Reinforcement Learning (MARL) methods and thereby show that the minimum price rule, in our setting, is generally difficult to achieve but far from impervious to tacit coordination among isolated decision-makers.
Our methodological choice, which relies on concepts defined in Section 2, is motivated by extensive literature that shows the effectiveness of combining game theory and reinforcement learning for modeling strategic interactions and optimizing behavior through dynamic adaptation among agents. We explore such literature in Section 3 and show that our methodological framework is ideal for exploring emergent behaviors and potential tacit collusion in public procurement markets. In Section 4, we provide the mathematical model for the MPMG and discuss the benefits of the Markov property in our setting. Section 5 presents the results of our computer experiments and concludes on their implications for tacit coordination in minimum price-driven auctions. Finally, Section 6 provides perspectives for future research and practical considerations.
2 Preliminaries
A Markov Decision Process (MDP) is defined by a tuple $(S,A,P,R,\gamma)$, where $S$ represents a finite set of states and $A$ is a finite set of actions. The state transition probability function $P(s^{\prime}|s,a)$ denotes the probability of moving from state $s$ to state $s^{\prime}$ given action $a$, while the reward function $R(s,a)$ specifies the immediate reward received after performing action $a$ in state $s$. The discount factor $\gamma$ determines the present value of future rewards, balancing immediate and long-term gains. A higher discount factor places greater emphasis on future rewards, encouraging the agent to consider long-term benefits, while a lower discount factor prioritizes immediate rewards. The goal in an MDP is to identify a policy $\pi$ that maximizes the expected cumulative reward. The key feature of an MDP is the Markov property, which asserts that future states depend only on the current state and the actions taken by the agents, not on the sequence of events that preceded it, ensuring that past states are irrelevant once the current state is known.
A Markov game is a generalization of a Markov Decision Process (MDP) to multiple interacting agents. The Markov property extends to this multi-agent setting, distinguishing Markov games from iterated games, which involve repeated play of a single-stage game with a static state representation based on past plays. Formally, a Markov game for $n$ players can be defined by
  • A set of states $\mathcal{S}$
  • For each player $i$, a set of actions $\mathcal{A}_{i}$
  • A transition function $T(s,a_{1},\dots,a_{n},s^{\prime})=\Pr(s^{\prime}\mid s,a_{1},\dots,a_{n})$ specifying the probability of moving from state $s$ to state $s^{\prime}$ given the actions $(a_{1},\dots,a_{n})$ taken by the players
  • For each player $i$, a reward function $R_{i}(s,a_{1},\dots,a_{n})$ specifying the immediate reward received by player $i$ after actions $(a_{1},\dots,a_{n})$ are taken in state $s$
Players aim to maximize their own total expected return, typically discounted over time. The solution concepts for Markov Games often involve finding Nash Equilibria, where no player can improve their payoff by unilaterally changing their strategy.
Multi-Agent Reinforcement Learning (MARL) extends single-agent reinforcement learning (RL) techniques to settings with multiple agents learning concurrently in a shared environment. In RL, an agent interacts with its environment to maximize cumulative rewards, where key components include agent, environment, state, action, reward, policy, and value function. The agent, acting as the decision-maker, responds to the environment that consists of external factors and situations represented as states. Actions are possible decisions the agent can take, evaluated by rewards as feedback on their success. The policy is the agent’s strategy for selecting actions based on the current state, while the value function estimates future rewards to determine the desirability of states or actions. In MARL, multiple agents simultaneously learn and update their policies, causing the environment to appear non-stationary from any single agent’s perspective. This dynamic leads to more complex interactions compared to single-agent RL, requiring robust techniques to address the challenges of Markov games and multiagent systems in general.
Multi-Armed Bandit (MAB) algorithms represent a simplified form of RL or MARL, focusing on the exploration-exploitation trade-off without considering future states, that is, without a model of the environment. The concept of "arms" in multi-armed bandits is analogous to the action space available to each agent. Each "arm" or action represents a distinct stochastic process with its own, a priori unknown, probability distribution of rewards.
Q-learning is an off-policy algorithm for RL and MARL problems where agents learn the value of actions directly by using a Q-value function to predict the expected utility of taking a given action in a given state. The Q-value function, denoted as $Q(s,a)$, represents the expected future rewards when an agent takes action $a$ in state $s$ and follows the optimal policy thereafter. The agent updates these values based on the rewards received. Off-policy means that the agent can learn the value of the optimal policy independently of the agent’s current actions, typically by using a replay buffer to store and reuse past experiences. This allows the agent to learn from data generated by different policies or even random actions, not just the current policy it is following. This class of methods is widely used for discrete action space problems because, in contrast to continuous problems, the state-action pair provides a clear and direct signal for learning.
Policy gradient methods focus on optimizing the policy directly by gradient ascent on expected rewards for RL or MARL problems. In MARL, policy gradient techniques, including the actor-critic framework, are used by each agent to adjust their policies in a way that considers the impact of other agents’ strategies, often involving complex inter-agent coordination and the handling of high-dimensional action spaces. The policy $\pi_{\theta}$ is updated by gradient ascent; namely
$\nabla J(\theta)=\mathbb{E}_{\pi_{\theta}}\left[\nabla_{\theta}\log\pi_{\theta}(a|s)Q^{\pi}(s,a)\right],$
where $\theta$ represents the parameters of the policy, and $Q^{\pi}(s,a)$ is the expected return following action $a$ in state $s$ under policy $\pi$.
3 Related Work
The interplay among MDPs, auction design, and multi-agent learning has gained significant attention in recent research. Although the literature addressing algorithmic collusion within this framework is relatively sparse, our study draws inspiration from the growing body of work on integrating Stackelberg game models within multi-agent learning frameworks, MDPs for auction designs, and tacit coordination in social dilemmas. By reviewing these interconnected areas, we aim to highlight the advancements in understanding strategic interactions, optimal bidding strategies, and tacit coordination among agents. This overview highlights the pivotal role of MDPs and related methodologies in shaping contemporary economic and strategic decision-making processes.
MDP and auction design. The MDP framework facilitates the exploration, understanding, and prediction of strategic interactions and behaviors within market designs. Examples include computing Nash equilibria in auction games [24, 25] and optimizing bidding strategies in combinatorial auctions [26, 27]. Although the Markov game framework is influential, bandit algorithms have also proven to be exceptionally suitable for modeling optimal bidding strategies in online auction designs [28, 29, 30, 31]. This is due to the simplicity and effectiveness of bandit algorithms, which operate in a stateless or single-state environment and make decisions independently of past actions without the need for an underlying state transition model.
MDP and Stackelberg game. Stackelberg models, in which players move sequentially rather than simultaneously and therefore were historically used to study hierarchical or sequential leadership, are being increasingly used in various applications within the Markov game and multiagent learning framework [32]. The Stackelberg game is a staple in the analysis of various economic and strategic situations where the ability to move first or commit to a decision before one’s competitors can confer a strategic advantage. This includes industries with clear market leaders who can influence the market environment [33]. For example, RL agents have been trained to learn the Stackelberg equilibrium in a Markov decision process representing economic mechanisms for assigning items to individualistic agents in sequential-pricing stages [34, 35, 36].
Multiagent learning, social dilemmas, and tacit coordination. In the exploration of tacit collusion and cooperative strategies within repeated game frameworks, a multitude of studies have contributed to a nuanced understanding of how agents can implicitly coordinate their actions under restricted communication conditions to optimize collective outcomes. The effectiveness of collusion in auction settings without explicit communication is a recurring theme [37], with similar phenomena observed in oligopoly models through Q-learning mechanisms where collusion emerges without direct interaction among firms [38, 39]. These findings are complemented by insights into multi-agent control using deep reinforcement learning, which further illustrates how complex cooperative behaviors can be engineered even in large-scale agent systems [40].
In scenarios where direct communication is either limited or non-existent, learning algorithms play a pivotal role in facilitating implicit cooperation and understanding competitive behaviors. For instance, learning methods that incorporate the anticipation of other agents’ learning processes have shown promise in fostering cooperative strategies that are robust against exploitation [41]. This is particularly evident in the Iterated Prisoner’s Dilemma, where strategies developed through evolutionary algorithms have been demonstrated to outperform traditional approaches [42]. Indeed, the contradictory nature of these games makes their study with RL agents quite challenging, as demonstrated in [43]. Moreover, the development of frameworks that enforce cooperation and resist collusion highlights the potential for creating stable cooperative environments in public goods games and other multi-agent settings [44, 45]. These strategies ensure that cooperation is not only a temporary state but a stable equilibrium, which is crucial for the long-term sustainability of collective actions in decentralized systems.
In this paper, we build on this literature and explore the potential for tacit coordination among isolated artificial agents in our social dilemma-based MPMG using multiagent learning.
4 The Minimum Price Markov Game
In public procurement, there is a tension between the pursuit of personal gain and collective welfare. Firms strive to maximize their individual profits, while the state aims to allocate public funds efficiently and maintain market competition and fairness. From the firms’ perspective, individually rational choices made without cooperation can lead to suboptimal outcomes for all participants, as maximal profits are only achievable through cooperative behavior. This conflict between individual and collective rationality is the core characteristic of social dilemmas. The minimum price rule, with its mathematical elegance, does not diminish this inherent conflict; rather, it preserves this tension fully, maintaining the balance between self-interest and group [^1] For clarity, here, the term ”group” refers to the group of firms, in contrast to the terms ”common” or ”collective”, which we employ to denote the whole economy. benefit.
It follows that a minimum price-ruled auction game formulation will fall in accordance with the principles of social dilemma games. Namely, it must display non-zero-sum interactions for which players can benefit from cooperating or suffer from a single defection. The dominant strategy (Nash equilibrium), being collectively irrational, is thus Pareto inefficient, meaning that the group optimal outcome aligns with a Pareto equilibrium. An efficient coordination scheme leads to the strategy profile associated with the profit-maximizing Pareto equilibrium. In other words, the role of coordination is to transition from the Nash to a Pareto equilibrium.
In this section, we provide a formulation for the MPMG, which we believe is a reasonable approximation of the typical real-world minimum price-ruled public procurement process. Our formulation is grounded in assumptions that capture the core essence of this process, creating a conducive environment to observe how artificial learners dynamically shape their strategies in such dilemma. The MPMG is founded on a static normal form game, the Minimum Price Game (MPG), which we will define first.
4.1 The Minimum Price Game: A Single-Stage Formulation
The MPG involves a public contractor, implicitly represented by the game’s rules and payoff structure, along with $n$ firms competing to win a contract in a minimum price auction, where the lowest bidder wins. Let $\mathcal{N}=\{1,\dots,n\}$ represent the set of firms in a specific industry. Each firm $i$ has a market power defined by the parameter $\beta_{i}\in(0,1)$, with $\sum_{i\in\mathcal{N}}\beta_{i}=1$, indicating that the average market strength, $\mu(\beta)$, is always $\frac{1}{n}$. This market power might represent the size of the firm or its market share, influencing its ability to reduce costs and leverage economies of scale. All firms share the same cost function, which determines their bids
$b_{i}=(1-\beta_{i})v\quad\forall i\in\mathcal{N},$ (1)
where $v$ is the value of the contract. The notion of strong and weak firms is relative to the distribution of $\beta$, but according to Equation (1), the closer $\beta$ is to 1, the lower the firm can bid, thus the stronger it is. Market heterogeneity is quantified by $\sigma(\beta)$, the standard deviation of the distribution of power parameter values. In a perfectly homogeneous market, $\sigma(\beta)=0$ and $\beta_{i}=\frac{1}{n}$ for all $i$. In such cases, the set of bids could display multiple minima if firms submit their prices directly following Equation (1). In other words, the homogeneous case is analogous to a game with no private values. This possibility arises from strategic considerations, though we have not yet fully outlined the space of strategies. We denote by $\Omega\subseteq\mathcal{N}$ the set of firms associated with the lowest bids.
Assumption 1 (Single Public Contractor).
All public contractors share the same goal.
In real-world scenarios, the set of public contractors in a given procurement market usually includes many entities. Despite this plurality, it is safe to assume that all public contractors work towards the same overarching goal. Assumption 1 allows us to reduce the set of public contractors to a singleton, a realistic simplification that will help in formulating an elegant model.
Assumption 2 (Single Stage Auction).
The bidding process unfolds in a single round.
Assumption 2 indicates that payoffs are determined in a single round of bid submissions. This necessitates that the payoff function accommodates the possibility of $\Omega$ including multiple firms. A simple solution for handling ties is to randomly select a winner, either by using the uniform distribution ($i^{*}\sim\mathcal{U}(\Omega)$) or by adding a stochastic perturbation to the bids ($b_{i}+\epsilon_{i}$). However, random selection would lead to a more complicated stochastic game without realistic justification. Furthermore, the process of random selection in ties forces $\Omega$ into a singleton, which may be overly restrictive. In practice, the likelihood of a tie in such a scenario is relatively low, and when it occurs, discrimination is not random.
Instead, we advocate for shared payoffs in the event of a tie. Sharing payoffs among the lowest bidders is justified for two reasons. First, even from a static perspective, this approach preserves and highlights the tension between individual and group rationality, which is the feature essential to our formulation and endeavor. Second, the MPG intends to provide a theoretical basis for further dynamic implementations. In dynamic settings, distributing the payoff among several bidders in a tie situation is analogous to random winner selection, as the cumulative payoffs for each bidder equalize over multiple auctions for a given repeated strategy profile.
Assumption 3 (Simultaneous Play).
All firms submit their bids simultaneously.
Although bids are, in practice, collected over a certain period, they remain sealed and confidential. This ensures that all bids are (virtually) submitted at the same time, thus justifying Assumption 3.
Assumption 4 (Common Value).
The contract value is common for all firms.
Assumption 4 simplifies the commonly used Bayesian game formulation of first-price auction to reflect the conditions of Bertrand competition [46], where firms compete on prices in a transparent market. This assumption is supported by the uniformity of contract requirements and regulatory costs across bidders, market transparency, and the homogeneity in the evaluation of contract values. These factors standardize the perceived value of the contract, thereby aligning all firms’ bids based on a common economic assessment rather than individual speculative valuations. Assumptions 1 to 4, combined with the simple one-dimensional minimum price rule, lead to a deterministic normal-form game formulation.
Strategies. All players are symmetric in their strategy set: for each firm $i$, $s_{i}=\{FP,CP\}$, where Fair Price (FP) is associated with bidding the fair price $b_{i}$ stemming from Equation (1), and Collusive Price (CP) with bidding a higher price. The collusive bid is defined by $\alpha\cdot b_{i}$, where $\tau\geq\alpha>1$, with the upper bound $\tau$ ensuring realistic bids. We denote the strategy profile for a given occurrence as the tuple $(s_{1},s_{2},\dots,s_{n})$. In the static setting, the specific bid values as well as the actual contract value are illustrative and not numerically significant since the payoffs are directly tied to the strategy profiles played.
Payoffs. Let $u_{i}(FP,k)$ be the payoff of firm $i$ if they bid the fair price while $k$ competitors bid the fair price, and $u_{i}(CP,k)$ the payoff of firm $i$ if they bid the collusive price while $k$ competitors bid the fair price. The total market power of lowest-bidding firms is
$\beta_{\Omega}=\sum_{j\in\Omega}\beta_{j}.$ (2)
The minimum price rule fosters a tension between group and individual rationality, and this dilemma must be represented in the payoff structure. The key idea is that coordination must be unanimous to achieve a collusive spoil, meaning that $u_{i}(CP,k=0)>u_{i}(FP,k=n-1)$. However, when $k=0$, that is, when all competitors bid the collusive price, $u_{i}(PF,k)>u_{i}(CP,k)$. Having in mind the fact that payoffs are shared in case of ties (following Assumption 2), and using Equation (2), we define the resulting payoff for player $i$ when they bid the fair price while $k=0$ as
$u_{i}(FP,k)=\frac{\beta_{i}}{\beta_{\Omega}}\cdot\alpha\cdot b_{i}.$ (3)
Conversely, when $k>0$, that is, if at least one competitor bid the fair price, $u_{i}(PF,k)>u_{i}(CP,k)$. Again following Assumption 2, and using Equation (2), we define the payoff for the player $i$ bidding the fair price while $k>0$ as
$u_{i}(FP,k)=\frac{\beta_{i}}{\beta_{\Omega}}\cdot b_{i}.$ (4)
In the static setting, actual payoff values $u_{i}(PF,k)$ when $k=0$ and $u_{i}(CP,k)$ when $k>0$ are not crucial as long as the payoff inequalities are respected.
Analysis. Equations (2), (3), and (4) define the payoff structure in the general heterogeneous case where the market power $\beta_{i}$ differs among firms. A perfectly homogeneous market is a special case where all players share the same market power, meaning $\beta_{i}=\frac{1}{n}$ for all $i\in\mathcal{N}$. In such scenarios, when collusion is achieved (i.e., player $i$ bids the collusive price and $k=0$), Equation (2) becomes $\beta_{\Omega}=1$ since $|\Omega|=n$, and hence Equation (3) becomes $u_{i}(CP,k)=\frac{1}{n}\cdot\alpha\cdot b_{i}$. Similarly, when player $i$ bids the fair price alongside at least one competitor (i.e., $k>0$), Equation (4) becomes $u_{i}(FP,k)=\frac{1}{(k+1)}\cdot b_{i}$ for all $i\in\Omega$.
Proposition 1.
The 2-player homogeneous Minimum Price Game (MPG), i.e., when both players share the same market power, as derived from Assumptions 1 to 4, is a 2-player classical Prisoner’s Dilemma.
Proof.
Suppose $n=2$, the game being composed of player $i$ and its opponent player $o$, both sharing the market equally ($\beta_{i}=\beta_{o}=0.5$). According to the payoff structure of the MPG, we can define each strategy profile for the 2-player case as follows:
  • $T$: Temptation payoff (bids the fair price while the other does not),
  • $R$: Reward payoff (both bid the collusive price),
  • $P$: Punishment payoff (both bid the fair price),
  • $S$: Sucker’s payoff (bids the collusive price while the other does not).
The 2-player homogeneous MPG is a classic Prisoner’s Dilemma as it respects the following three sufficient conditions:
  1. 1.
    $(FP,FP)$ is the dominant strategy profile.
    The minimum price rule inherently favors the lowest bidder. Thus, if player $o$ bids the fair price, the payoffs are
    $u_{i}(FP,FP)=\frac{1}{2}\cdot b_{i}=P\quad\text{and}\quad u_{i}(CP,FP)=0=S,\quad\text{therefore}\quad P>S.$
    In addition, since the players are symmetrical, $u_{i}=u_{o}$ for any strategy profile $(s_{i},s_{o})$. This means that bidding the fair price is the best response for both players, regardless of their opponent’s strategy. Thus, the strategy profile $(FP,FP)$ is the unique Nash equilibrium.
  2. 2.
    The strategy profile $(CP,CP)$ is Pareto dominant.
    A strategy profile is Pareto-efficient if no other strategy can make one player better off without making the other worse off. Since
    $u_{i}(CP,CP)=\frac{1}{2}\cdot\alpha\cdot b_{i}>u_{i}(FP,FP)=\frac{1}{2}\cdot b_{i}\quad\text{because}\quad\alpha>1,$
    $R>P$, meaning that $(CP,CP)$ is Pareto-efficient.
    A strategy profile is Pareto dominant if it is Pareto-efficient, and there is no other Pareto-efficient strategy profile that provides a higher payoff for at least one player without reducing the payoff for the other. The profiles $(CP,FP)$ and $(FP,CP)$ are also Pareto-efficient because the resulting bi-matrix game is symmetric, i.e,
    $u_{i}(CP,FP)=u_{o}(FP,CP)=S.$
    Since $R>P$ and $R>S$, the strategy profile $(CP,CP)$ is Pareto dominant.
  3. 3.
    The payoff structure respects the inequality $T>R>P>S$.
    We have already shown that $R>P>S$. Since
    $u_{i}(FP,CP)=b_{i}>u_{i}(CP,CP),$
    $T>R$, and therefore, $T>R>P>S$.
Corollary 1.
The n-player homogeneous MPG is a n-player prisoner’s dilemma.
Proof.
The 2-player MPG is a 2-player Prisoner’s Dilemma (Proposition 1). It is well established in game theory that the characteristics of the 2-player Prisoner’s Dilemma generalize to the $n$-player case. Therefore, the $n$-player homogeneous MPG also exhibits the same properties, making it an $n$-player Prisoner’s Dilemma. ∎
Therefore, the homogeneous MPG provides a simple yet profound framework for the study of strategic interactions among bidders in the context of minimum price-ruled auction markets. Mutual defection leads to suboptimal outcomes for all parties involved, but aligns with the prediction of market equilibrium. If firms were to cooperatively opt for higher prices, the group outcomes would be improved, highlighting the tension between competitive strategies and potential cooperative gains.
In the heterogeneous case, however, the market power $\beta$ differs among firms. This disparity leads to a variation in the payoffs. For a low-market power firm, bidding the fair price becomes more attractive because the payoff $u_{i}(PF,CP)$ can exceed the payoff $u_{i}(CP,CP)$ when all cooperate. The proportion of the reward $\frac{\beta_{i}}{\beta_{\Omega}}$ determines each firm’s payoff, potentially resulting in higher rewards for firms with larger market shares. Consequently, the strategy profile $(CP,CP)$ is no longer Pareto dominant because the incentive to bid the fair price is stronger for firms with lower market power, and the heterogeneous MPG is hence not a classical Prisoner’s Dilemma.
4.2 Markov Game Formulation
In contrast to iterated games, Markov games evolve over multiple stages, with states dependent on previous actions. This dynamic nature better captures the complexities of real-world markets where strategies evolve over time, especially in the context of behavioral emergence, which may heavily rely on self-strengthening trends. The MPMG extends the framework of the MPG by providing the advantage of the Markov property, which relies on a state space that encodes information about opponents’ past plays and characteristics.
Action space. The action space $\mathcal{A}=\{0,1\}$ is binary, reflecting the two strategies available to players, with bidding the fair price being represented by 00 and bidding the collusive price with $1$. This representation simplifies the computation of action and joint action frequencies. Let $a_{i}^{t}\in\mathcal{A}$ be the action taken by agent $i$ at time $t$, and $f^{t}\in\mathcal{A}^{n}$ be the joint action $(a_{1},\dots,a_{n})$ in auction $t$. The action and joint action frequencies up to iteration $t$ are then respectively given by
$\bar{a}_{i}^{t}=\frac{1}{t}\sum_{m=1}^{t}a_{i}^{m},$
and
$\bar{f}^{t}=\frac{1}{t}\sum_{m=1}^{t}f^{m}.$
Assumption 5 (Information availability).
All players have access to information about the game status and basic features of their opponents.
Assumption 5 is supported by the availability of auction data, which results from transparency laws. This assumption is crucial for constructing the state space, as it provides the signals necessary for players to make informed decisions during implementation.
State space. The choice of state space in a Markov game is often at the discretion of the implementer, depending on the specifics of the problem and the goals of the implementation. However, for our purposes, we define the MPMG with core variables that capture essential aspects of the game. Specifically, these sets include all players’ information on action frequencies $\mathbf{\bar{a}}^{t}=\{\bar{a}_{j}^{t}\}_{j\in\mathcal{N}}$, joint action frequencies $\mathbf{\bar{f}}^{t}=\{\bar{f}_{j}^{t}\}_{j\in\mathcal{N}}$, and power parameters $\mathbf{\beta}=\{\beta_{j}\}_{j\in\mathcal{N}}$. Therefore, the state space at any time $t$ is $\mathcal{S}=\{\mathbf{\bar{a}}^{t},\mathbf{\bar{f}}^{t},\mathbf{\beta}\}$, and is of size $2n+|\mathcal{A}|^{n}$. This minimal configuration for the state space aligns with Assumption 5. Furthermore, the power parameters also allow for the possibility of modeling a dynamic market in which firms could lose or gain market power.
Reward function. The payoff structure defined for the MPG translates to a reward function in the MPMG, emphasizing the inherent connection between the Markov game and its RL implementation. The reward function, together with the transition function, which defines how the game state evolves in response to actions, fundamentally shapes the dynamics and outcomes of the game. In fact, in the classical RL setting, the normalized reward function for firm $i$ at each iteration $t$ is expressed as:
$G_{i}^{t}=(1-\gamma)\sum_{m=0}^{t-1}\gamma^{m}r_{i}^{t-m},$
where $\gamma$ is the discount factor and $r_{i}^{t-m}$ is the reward received by firm $i$ at time $t-m$. The MPMG relies on the tension between Nash and Pareto dominant strategies inherent in the classical Prisoner’s Dilemma. However, as with the state space, the reward function can be adjusted for specific implementation purposes (e.g., reward shaping) as long as the fundamental dynamics are preserved. Following Assumption 4, we can further assume that the contract value $v$ remains equivalent as the iterations unfold, normalizing them to $1$. Equation (1) then becomes
$b_{i}=1-\beta_{i}.$
We do not make this assumption formal, as this normalization does not affect the generality of the model but only simplifies the calculations. The resulting bid $b_{i}$ can also be normalized to express a fair bid of $1$ and a collusive bid of $\alpha\cdot b_{i}=\alpha$, where $\tau\geq\alpha>1$. Considering these numerical adjustments, we get the immediate reward for player $i$ at any auction $t$ in the homogeneous case as
$r_{i}=\begin{cases}\frac{1}{(k+1)}&\text{if }a_{i}=0\\ \frac{\alpha}{n}&\text{if }a_{i}=1\text{ and }k=0\\ 0&\text{if }a_{i}=1\text{ and }k>0\end{cases}.$
Note that for simplicity, we dropped the subscript $t$. When firms are heterogeneous, the immediate rewards are weighted by their market power, as in Equation (3) and (4):
$r_{i}=\begin{cases}\frac{\beta_{i}}{\beta_{\Omega}}\cdot\frac{1}{(k+1)}&\text{if }a_{i}=0\\ \frac{\beta_{i}}{\beta_{\Omega}}\cdot\frac{\alpha}{n}&\text{if }a_{i}=1\text{ and }k=0\\ 0&\text{if }a_{i}=1\text{ and }k>0\end{cases},$
where $\beta_{\Omega}$ is the total market power of the lowest-bidding firms. Note that in the heterogeneous case, the set of lowest bidders $\Omega$ is likely to be a singleton, depending on how power parameters are distributed.
5 Computer Experiments
We created a MARL environment following the MPMG formulation described in Section 4.2. In this setup, the MPMG is implemented as an episodic task, where each episode is one step long and corresponds to a single auction denoted as $t\in\{1,\dots,T\}$. Implementing the MPMG as an episodic task with one-step episodes aligns with Assumption 2 and simplifies the learning process by making state transitions and cumulative rewards straightforward. Implementation details are available in Appendix B. For the experiments, we set $\alpha=1.3$, meaning the collusive bid is 30% higher than the fair bid, as per the notations in Section 4.2.
When considering solution methods, techniques explicitly designed to foster cooperation, such as shared joint action spaces [47] and algorithmic communication [48], as well as those implicitly encouraging cooperation, like reputation-based models [49] and status-quo loss [50], are not suitable. Inducing cooperation in MARL settings is well-studied, but using these approaches would simulate scenarios where firms deliberately bias their pricing algorithms towards cooperation. Although this raises interesting questions, it is beyond the scope of this study. Nevertheless, we provide insights on this topic in Section 5.4. Here, we evaluate the inherent properties of the MPMG, requiring unbiased experiments achieved by assuming isolated firms. If collusion occurs in the MPMG, it would stem from the game’s inherent nature. Therefore, we used basic yet established state-of-the-art MARL algorithms. Additionally, since the artificial firms are considered isolated and do not benefit from sharing mechanisms (such as a common policy or state-value approximator), they remain symmetric in their algorithmic structures, supporting their decision-making processes. In other words, they share the same level of rationality.
In this section, we present the results of our computer experiments on the MPMG using three foundational MARL approaches: Multi-Armed Bandits (MAB), Deep Q-learning, and Actor-Critic Policy Gradient. As discussed in Section 2, MAB focuses on balancing the exploration-exploitation trade-off, Deep Q-learning aims at value function approximation for action selection, and Actor-Critic Policy Gradient integrates both value estimation and policy optimization. These methodologies represent the core families of techniques in MARL, providing a comprehensive perspective on their effectiveness in various scenarios.
A total of seven algorithms—three bandit variants, two Q-learning variants, and two policy gradient algorithms—were trained on four MPMG configurations: the 2-player homogeneous, the 5-player homogeneous , the 2-player heterogeneous, and the 5-player heterogeneous MPMG. In the heterogeneous cases, we set $\sigma(\beta)$ to 0.5. As discussed in Section 4, $0\leq\sigma(\beta)\leq 1$ represents the standard deviation of the power parameters ($\beta$) distribution, with $\sigma(\beta)=0$ representing the homogeneous case. Each algorithm was trained on 100 auctions over 100 replications for statistical robustness. Results are presented in graphs and tables in this section. Additional material, such as graphs about algorithmic convergence and other relevant metrics, can be found in Appendix A, while Appendix B contains implementation details (e.g., hyperparameter values).
5.1 Multi-Armed Bandits
Our first implementation involves representing the MPMG as a bandit problem because MAB algorithms are lightweight and can be used independently of a Markov game formulation. Thus, our first implementation with bandits resembles an iterated MPG, as we do not utilize the Markov property. Implementing MAB as a first step before tackling model-based MARL methods is advantageous due to its focus on the exploration-exploitation trade-off. MAB provides a low-cost approach to eventually uncovering core strategies, serving as a proof-of-concept essential to evaluating the MPMG. In this study, we used three bandit algorithms: $\epsilon$-greedy, Thompson Sampling (TS), and Upper Confidence Bound (UCB).
The $\epsilon$-greedy bandit algorithm is a simple yet effective strategy for balancing exploration and exploitation in multi-armed bandit problems. At each step, the algorithm selects a random action with probability $\epsilon$ (exploration) and the action with the highest estimated reward with probability $1-\epsilon$ (exploitation). This approach ensures that the algorithm occasionally explores new actions to discover potentially better rewards while primarily exploiting the best-known actions to maximize cumulative rewards.
TS is a Bayesian approach to the multi-armed bandit problem that balances exploration and exploitation by maintaining a probability distribution for the expected reward of each arm. At each step, it samples from these distributions and selects the arm with the highest sampled value. Over time, this method efficiently narrows down the best-performing arm while still exploring other options.
The UCB algorithm addresses the exploration-exploitation trade-off by selecting the arm with the highest upper confidence bound of the estimated reward. It adds a confidence interval to the estimated reward of each arm that decreases as more data is collected, encouraging the exploration of less-tried arms initially and favoring exploitation as more information is gathered.
The performance of our bandit algorithms is displayed in Figure 5.1 in terms of training joint action frequencies, which represent the strategy profiles played during training. Since bandit agents are stateless, presenting the results of evaluation episodes is unnecessary, as their strategies would not evolve as the game unfolds. By evaluation, we mean making agents play with their learned and crystallized policies. Figure 5.1 (a) and (b) demonstrate that both $\epsilon$-greedy and TS agents have learned to play the Nash equilibrium strategy, which involves bidding the fair price. This outcome is expected, as bandit algorithms are designed to optimize immediate rewards and naturally gravitate towards strategies that consistently yield the highest payoff. In the context of the iterated MPG, defecting is the dominant strategy, leading to higher individual payoffs when both agents defect. However, Figure 5.1 (c) shows that UCB agents consistently managed to coordinate towards collusion, with their policies translating into a 60% frequency for $(CP,CP)$ on average. As expected, the frequencies of $(CP,FP)$ and $(FP,CP)$ decrease as agents exploit more, reaching marginal values, while the Nash equilibrium strategy ranks second with a 35% frequency on average after 100 auctions.
Figure 5.1: Training joint action frequencies of $\epsilon$-greedy (a), TS (b), and UCB (c) on the 2-player iterated homogeneous $(n=2,\sigma(\beta)=0)$ MPG over 100 replications of 100 iterations each. The colored bands show the confidence intervals over 100 replications.
(a) $\epsilon$-greedy
(b) Thompson Sampling
(c) UCB
Moreover, as expected, $\epsilon$-greedy and TS agents did not manage to learn how to collude in the $n$-player heterogeneous MPMG (see Appendix A), as the number of agents $n$ and the level of heterogeneity $\sigma(\beta)$ are known to hinder the collusive potential of a market as coordination gets harder. Consequently, we also expect UCB agents to learn to play the Nash more often as $n$ or $\sigma(\beta)$ increases. In fact, this is what we observe in Figure 5.2. We can see that market heterogeneity and the number of agents both negatively and independently impact the market’s collusive potential. Note that Figure 5.2 (b) does not display any legend for clarity since there are $2^{n}$ possible joint actions; however, the red and grey plots, respectively, represent the Pareto dominant strategy and the Nash, being consistent with Figure 5.2 (a). As we can see from Figure 5.2 (c), not only market heterogeneity and the number of agents force UCB agents to shift towards the Nash, but when both factors are set to inconvenient values, the uncertainty increases.
Figure 5.2: Training joint action frequencies drawn by UCB agents in the 2-player heterogeneous (a), the 5-player homogeneous (b), and the 5-player heterogeneous iterated MPG (c). The colored bands show the confidence intervals over 100 replications.
(a) $n=2$, $\sigma(\beta)=0.5$
(b) $n=5$, $\sigma(\beta)=0$
(c) $n=5$, $\sigma(\beta)=0.5$
Since a residual amount of collusion can occur during training regardless of the agents’ direction, we have set a threshold to determine what can be called collusion. This threshold is 30% of the maximal collusive spoil possible. TS and $\epsilon$-greedy agents have learned the Nash equilibrium 100% of the time, and they never achieved a significant amount of collusion. However, UCB agents, in regard to our threshold, managed to collude 100% of the time on the 2-player homogeneous iterated MPG. Table 1 reports the statistics related to collusion for UCB agents, including the collusion rate, which tells how many times UCB agents have achieved a collusive spoil that exceeds the threshold.
Table 1: UCB agents’ performance regarding collusion on 100 training auctions over 100 replications. The collusion rate is given as a percentage of the number of replications where collusion exceeded the threshold of 30% of the potential maximal collusive spoil. The average, standard deviation (STD), max, and min spoil are given as a percentage of the maximal collusive spoil achievable. These statistics are given for several market configurations.
$n=2$, $\sigma(\beta)=0$ $n=2$, $\sigma(\beta)=0.5$ $n=5$, $\sigma(\beta)=0$ $n=5$, $\sigma(\beta)=0.5$
Collusion Rate 100 4 40 0
Average Spoil 64.5 9 25 0
STD Spoil 1.5 8.8 11.6 2.6
Max Spoil 66 59 57 15
Min Spoil 63 0 17 0
Table 1 shows that market heterogeneity impedes collusive potential more than the number of agents, as UCB agents have achieved collusion 40% of the time in the 5-player homogeneous MPG and only 4% of the time in the 2-player heterogeneous MPG. Moreover, the average collusive spoil is much higher in the 5-player homogeneous case than in the 2-player heterogeneous case. When both market heterogeneity and the number of agents increase, collusion is fully tamed, as the average collusive spoil in this case is 0% of the maximal spoil achievable.
5.2 Deep Q-learning
Assuming isolated firms, each agent can treat competitors as part of the environment and learn a Q-value function independently, considering the impact of other agents’ actions as part of the state transitions. This leads to the following representation for a given state:
$Q_{i}(s,a_{i})=\mathbb{E}\left[R_{i}\mid s,a_{i}\right],$ (5)
where $s$ encodes information about all players. When opponent modeling is used, opponents’ behaviors are explicitly incorporated into the Q-value function. The enhanced Q-value function, denoted as $Q(s,a,s_{o},a_{o})$, incorporates the opponent’s state $s_{o}$ and action $a_{o}$ directly into the learning process. For example, with 2 agents
$Q(s,a,s_{o},a_{o})=\mathbb{E}[r+\gamma\max_{a^{\prime}}Q(s^{\prime},a^{\prime},s^{\prime}_{o},a^{\prime}_{o})|s,a,s_{o},a_{o}],$ (6)
where $s$ and $s_{o}$ are the states of the agent and the opponent, respectively, $a$ and $a_{o}$ are their corresponding actions, $r$ is the reward, $\gamma$ is the discount factor, and $a^{\prime}$ and $a^{\prime}_{o}$ are the next actions of the agent and the opponent.
In our experiments, we used the Double Deep Q-Network (D3QN) [51] in both the versions depicted by Equation (5) and (6). We refer to the former case as naive D3QN, and the latter as D3QN with opponent modeling (D3QN-OM). Figure 5.3 shows that the MPMG is robust to tacit coordination among rational agents. The naive D3QN may be naive in its structure compared to the D3QN-OM, but it still encodes a higher level of rationality than bandit algorithms that do not use state representation. From Figure 5.3 (a), we can see that D3QN agents converge with good certainty towards the Nash equilibrium. Indeed, D3QN agents learn to assign higher Q-values to action 0 as the game unfolds, as described by Figure 5.3 (b) and (c). It follows without surprise that the naive D3QN agents have learned to play along the same pattern when either the number of agents or market heterogeneity increases (see Appendix A).
Figure 5.3: Joint action frequencies of the naive D3QN during evaluation (a), training Q-values for action 0 (b), and training Q-values for action 1 (c) in the 2-player homogeneous $(n=2,\sigma(\beta)=0)$ MPMG. The colored bands show the confidence intervals over 100 replications.
(a) Joint action frequencies on evaluation episodes
(b) Q-values for action Fair Price during training
(c) Q-values for action Collusive Price during training
It turns out that the impact of opponent modeling on Q-agents’ ability to discover profit-maximizing strategies in the 2-player homogeneous MPMG is small. Indeed, D3QN-OM agents achieve some amount of collusive spoil, which is more than the naive D3QN agents did. However, as we can observe from Figure 5.4 (a), D3QN-OM agents seem uncertain. Indeed, the Nash equilibrium has an average rate of near $0.5$, and its associated confidence interval is rather wide. Nevertheless, D3QN-OM agents, on average, converge towards the Nash equilibrium, as shown in Figure 5.4 (b) and (c). Additionally, as with the naive D3QN, we observe no surprises regarding the agent behavior in the general $n$-player heterogeneous setting (see Appendix A).
Figure 5.4: Joint action frequencies of the D3QN-OM agents during evaluation (a), training Q-values for action 0 (b), and training Q-values for action 1 (c) in the 2-player homogeneous $(n=2,\sigma(\beta)=0)$ MPMG. The colored bands show the confidence intervals over 100 replications.
(a) Joint action frequencies on evaluation episodes
(b) Q-values for action Fair Price during training
(c) Q-values for action Collusive Price during training
Despite the fact that D3QN agents never passed the collusive threshold, their collusive potential exists, and the D3QN-OM method is somewhat an improvement from the naive D3QN as we can see in Table 2.
Table 2: D3QN-OM agents’ performance regarding collusion on 100 training auctions over 100 replications. The collusion rate is given as a percentage of the number of replications where collusion exceeded the collusive threshold of 30% of the potential maximal collusive spoil. The average, standard deviation (STD), max, and min spoil are given as a percentage of the maximal collusive spoil achievable. These statistics are given for several market configurations.
$n=2$, $\sigma(\beta)=0$ $n=2$, $\sigma(\beta)=0.5$ $n=5$, $\sigma(\beta)=0$ $n=5$, $\sigma(\beta)=0.5$
Collusion Rate 0 0 0 0
Average Spoil 17.6 17.5 0 0
STD Spoil 11.5 11.1 0 0
Max Spoil 30 33 8 0
Min Spoil 0 0 0 0
5.3 Actor-Critic Policy Gradient
In the last experiments, we used Multi-Agent Proximal Policy Optimization (MAPPO), a policy gradient algorithm where the policy represents the probability distribution over bidding either the fair price or a collusive price. MAPPO employs the actor-critic method, in which a critic evaluates the actions taken by the actor by estimating the value function—representing expected returns from the current state following the current policy. The actors are updated based on the gradients of the expected reward, as assessed by the critic. The goal of the actors is to refine the policy using feedback on their performance from the critic.
Figure 5.5: Joint action frequencies during evaluation (a), training policy for action 0 (b), and training policy for action 1 (c) of MAPPO agents in the 2-player homogeneous $(n=2,\sigma(\beta)=0)$ MPMG. The colored bands show the confidence intervals over 100 replications.
(a) Joint action frequencies on evaluation
(b) Policy for action Fair Price during training
(c) Policy for action Collusive Price during training
We found that MAPPO agents always learn to play the Nash, regardless of market settings. As suggested by Figure 6.3, which shows the MAPPO results on the 2-player homogeneous MPMG, tacit coordination does not occur in the $n$-player heterogeneous setting (see Appendix A). Figure 6.3 (b) and (c) show that MAPPO agents neatly and consistently converge towards the Nash play.
5.4 Conclusion
The results of our experiments underscore the varying effectiveness of different MARL approaches in the MPMG. From our observations in Section 5, three key insights emerge.
First, there appears to be no positive correlation between the level of complexity of our AI agents and the emergence of tacit coordination. Interestingly, only the UCB agents, utilizing a relatively simple method, consistently achieved substantial collusive spoils while demonstrating robustness in non-favorable market conditions (see Table 1). This may seem counterintuitive, as greater algorithmic complexity and knowledge are typically expected to provide more perspective. However, it is important to note that the rational strategy is to play the Nash equilibrium, as humans tend to do in similar settings [52]. Therefore, we can draw the following partial conclusion: assuming the MPMG reasonably approximates minimum price-ruled auctions, the minimum price rule is resilient to non-engineered tacit coordination among rational actors, though paradoxically, collusion can occur in the most naive settings.
Second, the emergence of tacit coordination does not appear to rely heavily on self-reinforcing trends. The state representation, evolving linearly as the game progresses, seems to favor the Nash equilibrium. This can be attributed to the fact that achieving the profit-maximizing Pareto equilibrium requires a fragile and precise set of conditions if it depends on self-reinforcing trends. In practice, players will always respond to cues of defection from their opponents, and since agents are symmetrical (at least in this study), no player is likely to make the first move. Coordination could theoretically be triggered by sustained mutual exploration of defection, which is even less probable when agents are fully isolated, as is the case here. The question of whether tacit coordination arises from self-reinforcing empirically-based beliefs or from planned behavior supported by encouraging beliefs about opponents remains open.
Third, the MPMG exhibits properties consistent with Jean Tirole’s seminal work on market collusion potential [53]. Specifically, as the number of firms increases, coordination becomes more challenging because the value of $u_{i}(CP,k=0)$ decreases, undermining the intrinsic value of the Pareto dominant strategy and thus reducing the likelihood of successful collusion. Additionally, market heterogeneity, characterized by varying market power or market share among firms, negatively impacts the incentive to coordinate. This supports the idea that the MPMG provides a simple yet realistic approximation of market dynamics.
6 Discussion
Our results align with specialized studies that reveal the challenge of achieving algorithmic cooperation, whether tacit or explicit [54]. However, they do not contradict the legal literature, where scholars often assert that algorithmic collusion is relatively straightforward to achieve [55, 56, 4, 57]. This study offers a framework and benchmark for tacit collusion and helps explain the contradictory reports in the literature by shedding light on the opaque nature of algorithmic tacit coordination.
Nonetheless, it is essential to add nuance to achieve a comprehensive understanding of this complex issue. While legal concerns stem from tangible evidence found in digital open markets and auctions (see Section 1), algorithmic implicit coordination has been shown to be impeded in the specific context of unbiased social dilemmas, as discussed in Section 3. Additionally, the potential for algorithmic and tacit collusion in transparent markets (including public markets) will always exist. Complex cooperative behaviors can be engineered in virtually any large-scale agent system [40], even if such implementations remain challenging. For instance, cooperation in competitive settings can be elusive even when explicit communication is permitted [58, 59]. What is certain, at least from our perspective, is that further research is required to establish the precise theoretical and algorithmic conditions, as well as the practical considerations, regarding algorithmic collusion in public procurement markets.
Extending the MPMG. Future research should aim to extend the MPMG in both quantitative and qualitative dimensions. Quantitatively, this could involve creating environments that more closely mimic actual simulations, using approaches like empirical games [60]. Qualitatively, it could involve relaxing certain assumptions, such as the common value assumption (Assumption 4), particularly in highly heterogeneous market scenarios. Ultimately, developing a realistic, full-scale simulation of such a system would also necessitate a continuous formulation for the action space and a modular environment allowing for the implementation of various supplier selection models, thus generalizing the MPMG to the Public Procurement Markov Game (PPMG).
Additionally, the public contractor could be modeled as an active participant in the environment, represented by an adaptive learning agent. This scenario has been explored by [34], where reinforcement learning was used in the context of Stackelberg games as an algorithmic defense against algorithmic pricing for e-commerce platforms. Practically, these perspectives could shape the future of more robust E-procurement designs, and theoretically, they offer a computer-based experimental framework that complements laboratory-based behavioral studies.
AI rationality and cyber cartels. In an AI-governed world, the concept of rationality takes on new dimensions, particularly in economic interactions and market behavior. The experiments in this study operate under a set of restrictions to maintain an environment unbiased towards cooperation. However, as mentioned in Section 5, market players could align on algorithmic pricing using centralized learning and execution (i.e., a shared joint action space). This raises the question: could tacit collusion, emerging from a deliberately biased algorithmic structure, be easily detectable?
The potential existence of cyber cartels introduces new research opportunities into how firms might evade detection and punishment by law enforcers through planned algorithmic collusion. Such a consortium could gain advantages in both efficiency and stealth. Explicit coding for collusion would typically require real-life coordination among market players, potentially leaving traces and being susceptible to algorithmic auditing by legal authorities. However, implicit algorithmic coordination can either be programmed or accidental, creating a grey zone that complicates regulatory efforts. It is also worth noting that both digitized and non-digitized markets can be affected, as the only requirement for algorithmic pricing is the availability of data. Distinguishing between intentional and unintentional collusion requires sophisticated analytical tools and a deep understanding of AI behavior to achieve cooperation, highlighting the need for specialized literature and expertise to foster regulatory adaptation.
Competing Interests
The authors declare no competing interests.
Author Contributions
All authors contributed to the study’s conception and design. Computer experiments were designed and performed by Igor Sadoune. The first draft of the manuscript was written by Igor Sadoune and all authors commented on previous versions of the manuscript. All authors read and approved the final manuscript.
Appendix A: Additional Figures
Figure 6.1: Loss values of the dueling DQN during training of the naive D3QN on the 2-player homogenous MPMG (a) and 5-player heterogeneous MPMG (b).
(a) DQN losses on the 2-player homogenous MPMG
(b) DQN losses on the 5-player heterogeneous MPMG
Figure 6.2: Loss values of the dueling DQN during training the D3QN with opponent modeling on the 2-player homogenous MPMG (a) and 5-player heterogeneous MPMG (b).
(a) DQN losses on the 2-player homogenous MPMG
(b) DQN losses on the 5-player heterogeneous MPMG
Figure 6.3: Actor network loss values during training of MAPPO on the 2-player homogenous MPMG (a) and 5-player heterogeneous MPMG (b). Simalarly, graphs (c) and (d) shows the associated critic network loss values for the 2-player homogeneous and 5-player heterogeneous case.
(a) Actor network losses on the 2-player homogenous MPMG
(b) Actor network losses on the 5-player heterogeneous MPMG
(c) Critic network losses on the 2-player homogenous MPMG
(d) Critic network losses on the 5-player heterogeneous MPMG
Appendix B: Implementation Details
In this section, we expose in detail our implementation of the experiments presented in this manuscript. The code associated to this study is written in Python, and can be found at https://github.com/IgorSadoune/minimum-price-game.
E-Greedy Method. Our EpsilonGreedyMAB class utilizes an $\epsilon$-greedy strategy with the following hyperparameters:
  • n_arms: Number of possible actions (arms) for the bandit.
  • epsilon: Exploration rate, set to 0.3 in the initialize_agents function.
The agent selects a random action with probability $\epsilon$ and the best known action with probability $1-\epsilon$. The action values are updated using an incremental average formula
$Q_{t+1}(a)=Q_{t}(a)+\frac{1}{N(a)}(R_{t}-Q_{t}(a))$
where $Q_{t}(a)$ is the estimated value of action $a$ at time $t$, $N(a)$ is the number of times action $a$ has been selected, and $R_{t}$ is the reward received at time $t$.
Upper Confidence Bound. Our UCBMAB class implements the UCB algorithm with the following hyperparameters:
  • n_arms: Number of possible actions (arms) for the bandit.
The agent selects actions based on confidence bounds calculated as:
$\text{UCB}_{t}(a)=Q_{t}(a)+\sqrt{\frac{2\log t}{N(a)}}$
where $Q_{t}(a)$ is the estimated value of action $a$ at time $t$, $N(a)$ is the number of times action $a$ has been selected, and $t$ is the total number of pulls. Action values are updated using the incremental average formula mentioned above.
Thompson Sampling. Our ThompsonSamplingMAB class uses a Bayesian approach to action selection with the following hyperparameters:
  • n_arms: Number of possible actions (arms) for the bandit.
  • $\alpha$: Beta distribution parameter
  • $\beta$: Beta distribution parameter
The agent maintains Beta distributions for each action, initialized with $\alpha=1$ and $\beta=1$ for each arm. Actions are selected based on samples drawn from these Beta distributions. The Beta distribution parameters are updated based on the received binary rewards as follows:
$\alpha_{\text{new}}$ $=\alpha_{\text{old}}+\text{reward}$
$\beta_{\text{new}}$ $=\beta_{\text{old}}+1-\text{reward}$
This method effectively balances exploration and exploitation by sampling from the probability distributions of each action’s expected reward.
Deep Dueling DQN (naive D3QN). The DuelingDQN class implements a Dueling Deep Q-Network with the following architecture:
  • feature_layer: A fully connected layer with 128 units followed by a ReLU activation function.
  • value_stream: A fully connected layer with 128 units followed by a ReLU activation function, and another fully connected layer with 1 unit representing the state value.
  • advantage_stream: A fully connected layer with 128 units followed by a ReLU activation function, and another fully connected layer with action_dim units representing the advantage values for each action.
The Q-values are computed as:
$Q(s,a)=V(s)+\left(A(s,a)-\frac{1}{|\mathcal{A}|}\sum_{a^{\prime}}A(s,a^{\prime})\right)$
where $V(s)$ is the state value, $A(s,a)$ is the advantage of action $a$ in state $s$, and $\mathcal{A}$ is the set of all possible actions.
The D3QNAgent class implements a Double Dueling Deep Q-Network agent with the following hyperparameters:
  • state_dim: Dimensionality of the state space.
  • action_dim: Dimensionality of the action space.
  • lr: Learning rate, set to $1\times 10^{-3}$.
  • gamma: Discount factor, set to 0.99.
  • epsilon: Initial exploration rate, set to 1.0.
  • epsilon_decay: Decay rate of epsilon, set to 0.995.
  • epsilon_min: Minimum exploration rate, set to 0.01.
  • buffer_size: Size of the replay buffer, set to 10,000.
The agent selects actions using an $\epsilon$-greedy strategy:
  • With probability $\epsilon$, select a random action.
  • With probability $1-\epsilon$, select the action with the highest Q-value:
    $a_{t}=\arg\max_{a}Q(s_{t},a)$
The agent stores experiences in a replay buffer and updates the Q-network by sampling mini-batches of experiences. The Q-values are updated using the Bellman equation:
$Q(s_{t},a_{t})=r_{t}+\gamma\max_{a^{\prime}}Q^{\prime}(s_{t+1},a^{\prime})$
where $Q^{\prime}$ is the target Q-network. The loss is computed using Mean Squared Error (MSE):
$\mathcal{L}=\frac{1}{N}\sum_{i=1}^{N}\left(Q(s_{t}^{i},a_{t}^{i})-y^{i}\right)^{2}$
where $y^{i}=r_{t}^{i}+\gamma\max_{a^{\prime}}Q^{\prime}(s_{t+1}^{i},a^{\prime})$. The target network is updated periodically by copying the weights from the Q-network.
Dueling DQN with opponent modeling (D3QN-OM). Our D3QN-OM shares the same DuelingDQN class with the naive D3QN as MLP representation. However, the OpponentModel class models the behavior of opponents with the following details:
  • state_dim: Dimensionality of the state space.
  • action_dim: Dimensionality of the action space.
  • model: A dictionary that maps states to action distributions.
The opponent model is updated based on observed actions, using the same D3QNAgent class, except for an instance of OpponentModel for opponent modeling. The opponent model is updated based on observed actions as follows:
  1. 1.
    If the current state is not already in the model, initialize the action distribution for this state as an array of zeros.
  2. 2.
    Increment the count for the observed action in the action distribution array corresponding to the current state.
The agent selects actions using an $\epsilon$-greedy strategy, but when opponent modeling is enabled, Q-values are adjusted based on opponent action distributions
$Q(s_{t},a)\times\text{opponent action distribution}$
The D3QN-OM agent interacts with stored experiences in a replay buffer as the same way the naive D3QN does.
Multiagent Proximal Policy Optimization (MAPPO). The MLPNetwork class implements a multi-layer perceptron network with the following architecture:
  • state_dim: Dimensionality of the state space.
  • action_dim: Dimensionality of the action space.
  • layers: A fully connected network with two hidden layers, each with 128 units followed by a ReLU activation function.
Our MAPPOAgent class implements a MAPPO agent with the following hyperparameters:
  • state_dim: Dimensionality of the state space.
  • action_dim: Dimensionality of the action space.
  • lr: Learning rate, set to $1\times 10^{-3}$.
  • gamma: Discount factor, set to 0.99.
  • clip_param: Clipping parameter for PPO, set to 0.2.
  • c1: Coefficient for value function loss, set to 0.5.
  • c2: Coefficient for entropy bonus, set to 0.01.
  • epsilon: Initial exploration rate, set to 1.0.
  • epsilon_decay: Decay rate of epsilon, set to 0.995.
  • epsilon_min: Minimum exploration rate, set to 0.01.
The agent selects actions using an $\epsilon$-greedy strategy:
  • With probability $\epsilon$, select a random action.
  • With probability $1-\epsilon$, select the action with the highest policy probability:
    $a_{t}=\arg\max_{a}\pi(a|s_{t})$
  • The exploration rate $\epsilon$ is decayed over time:
[…]
The agent stores experiences in a memory buffer for training:
  • Store experiences as tuples of (state, action, reward, next_state, done).
The Generalized Advantage Estimation (GAE) is computed using the following formula:
$\delta_{t}=r_{t}+\gamma V(s_{t+1})-V(s_{t})$
$\hat{A}_{t}=\sum_{l=0}^{\infty}(\gamma\lambda)^{l}\delta_{t+l}$
and the policy and value are updated as
  • Policy Update (Actor):
    $\text{ratio}=\frac{\pi(a_{t}|s_{t})}{\pi_{\text{old}}(a_{t}|s_{t})}$
    $\mathcal{L}^{\text{CLIP}}(\theta)=\mathbb{E}_{t}\left[\min(\text{ratio}\cdot\hat{A}_{t},\text{clip}(\text{ratio},1-\epsilon,1+\epsilon)\cdot\hat{A}_{t})\right]$
  • Value Update (Critic):
    $\mathcal{L}^{\text{VF}}=\left(\hat{R}_{t}-V(s_{t})\right)^{2}$
  • Total Loss:
    $\mathcal{L}=\mathcal{L}^{\text{CLIP}}+c_{1}\mathcal{L}^{\text{VF}}-c_{2}\mathcal{H}$
    where $\mathcal{H}$ is the entropy bonus.